-1/9(x-27)+1/3(x+3)=x-3
Please help solve the equation.

Answer by stanbon(75887)

-1/9(x-27)+1/3(x+3)=x-3
Multiply thru by -9 to get rid of the fractions:
(x-27) + -3(x+3) = -9(x-3)
---------------------
Simplify:
x - 27 - 3x - 9 = -9x + 27
---------------
Rearrange:
7x = 63
x = 9
============
Cheers,
Stan H.